# Modeling the One-to-Many Property in Open-Domain Dialogue with LLMs

### v1

@@ 6.1 Framework Evaluation
[…]
@@
It is also important to note that fine-tuning the ODRP model with hard negative samples leads to a noticeable improvement in the diversity and coherence of the selected responses across all LLMs. $ODRP$HN outperforms $ODRP$ on all automatic metrics and achieves a higher Win rate and lower Loss rate in human evaluation. The effectiveness of the ODRP model is particularly evident in the case of TinyLlama, where there is substantial variability in the quality of responses generated during MRG. Generally, we observe that the ODRP model excels at identifying and prioritizing higher-quality responses, resulting in more engaging and meaningful exchanges, even when the initial set of responses exhibits significant variability. This leads to improvements of up to 90% in response diversity and contextual coherence.
[…]
{tblr}
cell12 = c, cell13 = c, cell14 = c, cell15 = c, cell22 = c, cell23 = c, cell24 = c, cell25 = c, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell42 = c, cell43 = c, cell44 = c, cell45 = c, cell52 = c, cell53 = c, cell54 = c, cell55 = c, cell62 = c, cell63 = c, cell64 = c, cell65 = c, cell72 = c, cell73 = c, cell74 = c, cell75 = c, cell82 = c, cell83 = c, cell84 = c, cell85 = c, cell92 = c, cell93 = c, cell94 = c, cell95 = c, cell102 = c, cell103 = c, cell104 = c, cell105 = c, cell112 = c, cell113 = c, cell114 = c, cell115 = c, cell122 = c, cell123 = c, cell124 = c, cell125 = c, cell132 = c, cell133 = c, cell134 = c, cell135 = c, cell142 = c, cell143 = c, cell144 = c, cell145 = c, cell152 = c, cell153 = c, cell154 = c, cell155 = c, cell162 = c, cell163 = c, cell164 = c, cell165 = c, cell172 = c, cell173 = c, cell174 = c, cell175 = c, cell182 = c, cell183 = c, cell184 = c, cell185 = c, cell192 = c, cell193 = c, cell194 = c, cell195 = c, hline1-2,8,14,20,22 = -, rowsep=1.0pt & Dist-1 Dist-2 UE UniEval
$TinyLlama$ 0.16*/0.18* 0.51*/0.61* 0.21*/0.13* 0.74*/0.64*
- $rand$ 0.24*/0.20* 0.75*/0.70* 0.24/0.13* 0.76*/0.65*
- $cls$ 0.22*/0.25* 0.76*/0.74* 0.23/0.18* 0.78*/0.66*
- $pref$ 0.25*/0.24* 0.73*/0.75* 0.24/0.18* 0.77*/0.70*
- $ODRP$ 0.28*/0.29 0.77/0.798 0.27/0.22* 0.81/0.72*
- $ODRP$HN 0.31/0.31 0.79/0.82 0.30/0.26 0.83/0.76
$Llama2\text{-}7b$ 0.20*/0.22* 0.61*/0.69* 0.24*/0.21 0.83/0.72
- $rand$  0.23*/0.30* 0.77*/0.78* 0.22*/0.19* 0.81*/0.69*
- $cls$ 0.30*/0.27* 0.79*/0.75* 0.23*/0.18* 0.80*/0.72
- $pref$ 0.28/0.29* 0.77*/0.78* 0.24*/0.22 0.83/0.71*
- $ODRP$ 0.33/0.35 0.83/0.84 0.26/0.22 0.83/0.73
- $ODRP$HN 0.35/0.36 0.83/0.85 0.29/0.24 0.85/0.73
$Llama2\text{-}13b$ 0.21*/0.23* 0.65*/0.72* 0.26*/0.24* 0.85/0.77*
- $rand$  0.24*/0.28* 0.77*/0.76* 0.25*/0.24* 0.80*/0.72*
- $cls$ 0.30*/0.31* 0.80*/0.76* 0.29*/0.25 0.83*/0.77*
- $pref$ 0.31/0.30* 0.79*/0.78 0.26*/0.29 0.82*/0.79
- $ODRP$ 0.33/0.34 0.85/0.79 0.32/0.30 0.85/0.81
- $ODRP$HN 0.33/0.35 0.84/0.82 0.33/0.32 0.87/0.82
$Llama2\text{-}70b$ 0.31/0.32 0.72/0.80 0.28/0.26 0.86/0.79
$gpt\text{-}3.5\text{-}turbo$ 0.36/0.33 0.75/0.82 0.31/0.30 0.88/0.81
{tblr}
column3 = c, column4 = c, column5 = c, cell21 = r=7, cell91 = r=7, cell161 = r=7, hline1-2,9,16,23 = -, rowsep=1.0pt & Win Tie Loss
TinyLlama $ODRP$HN vs. $TinyLlama$ 85 9 6
$ODRP$HN vs. $rand$ 76 16 18
$ODRP$HN vs. $cls$ 60 29 11
$ODRP$HN vs. $pref$ 57 20 23
$ODRP$HN vs. $ODRP$ 49 33 18
$ODRP$HN vs. $Llama2\text{-}70b$ 30 35 35
$ODRP$HN vs. $gpt\text{-}3.5\text{-}turbo$ 26 44 30
Llama2-7b $ODRP$HN vs. $Llama2\text{-}7b$ 74 18 8
$ODRP$HN vs. $rand$ 58 25 17
$ODRP$HN vs. $cls$ 50 29 21
$ODRP$HN vs. $pref$ 47 27 23
$ODRP$HN vs. $ODRP$ 46 30 24
$ODRP$HN vs. $Llama2\text{-}70b$ 32 41 27
$ODRP$HN vs. $gpt\text{-}3.5\text{-}turbo$ 28 48 24
Llama2-13b $ODRP$HN vs. $Llama2\text{-}13b$ 50 33 17
$ODRP$HN vs. $rand$ 51 24 25
$ODRP$HN vs. $cls$ 44 34 22
$ODRP$HN vs. $pref$ 42 30 28
$ODRP$HN vs. $ODRP$ 41 32 27
$ODRP$HN vs. $Llama2\text{-}70b$ 38 39 23
$ODRP$HN vs. $gpt\text{-}3.5\text{-}turbo$ 37 40 23
@@ 6.2 MRG Evaluation
Comparison with MI Response sets generated via MI tend to be semantically similar despite relatively high lexical diversity, as shown by low inter-response semantic scores and comparably higher lexical diversity scores in both automatic and human evaluations. This is likely due to the deterministic nature of logits during inference. Although sampling strategies (temperature scaling Guo et al. (2017) or nucleus sampling Holtzman et al. (2020)) introduce stochasticity in decoding, generated logits remain deterministic, limiting semantic variation unless randomness is significantly increased, which could reduce contextual coherence.
@@ 7 Related Work
Prior work adopting a two-stage approach for open-domain dialogue typically involves generating multiple responses either through conditional generation based on pre-specified dialogue acts Sakaeda and Kawahara (2022) or by pooling outputs from variational and retrieval-based systems Ruan et al. (2020); of Physics and Technology (2021). However, these studies often focus on evaluating only the final selected response, without considering the diversity or contextual coherence of the entire set of generated responses. In contrast, our approach evaluates and optimizes the quality of the full set of responses, thereby enhancing the overall quality of the final selected response. Additionally, many of these methods have been applied to smaller language models, whereas to the best of our knowledge, our work is the first to introduce a two-stage generation framework LLMs. Other two-stage approaches broadly entail first generating a candidate response and instantiating it as the final response Li et al. (2023), or generating a response in the first stage and further conditioning and refining the response in the second stage Qian et al. (2024); Shi and Song (2023).
Regarding response selection, prior work has primarily concentrated on narrow criteria such as engagement Sakaeda and Kawahara (2022), topical relevance Ruan et al. (2020); Yuan et al. (2024). Standard retrieval-based systems, in contrast, prioritize contextual coherence Tao et al. (2021); Su et al. (2024). In our framework, we prioritize human preferences, considering a broader range of factors such as harmlessness and helpfulness, which are critical aspects for ensuring the real-world utility of response generation systems.

### v2

@@ 6.1 Framework Evaluation
Based on the results obtained, it is clear that the responses selected by $ODRP$ and ODRPODRPHN consistently outperform all other approaches, including randrand, clscls, and prefpref, in terms of both diversity and contextual coherence. Both $ODRP$ and $ODRP\textsubscript{HN}$ generally achieve statistically significantly higher Distinct and UE/UniEval scores than the baseline methods. Moreover, in human evaluation, they show a greater proportion of wins and a lower proportion of losses compared to other baselines. Qualitatively, we observe that responses selected by $ODRP$ and ODRPODRPHN do more than just acknowledge the previous utterance; they often provide additional enriching information that enhances the overall dialogue. Furthermore, a significant portion of these selected responses include queries directed at the other interlocutor, actively encouraging further interaction.
@@
It is also important to note that fine-tuning the ODRP model with hard negative samples leads to a noticeable improvement in the diversity and coherence of the selected responses across all LLMs. ODRPODRPHN outperforms $ODRP$ on all automatic metrics and achieves a higher Win rate and lower Loss rate in human evaluation. The effectiveness of the ODRP model is particularly evident in the case of TinyLlama, where there is substantial variability in the quality of responses generated during MRG. Generally, we observe that the ODRP model excels at identifying and prioritizing higher-quality responses, resulting in more engaging and meaningful exchanges, even when the initial set of responses exhibits significant variability. This leads to improvements of up to 90% in response diversity and contextual coherence.
Comparison with Larger LLMs In addition, we evaluated larger LLMs, such as Llama2-70b and gpt-3.5-turbo, using the zero-shot variant of the FS prompt (Figure 3). Our findings reveal that after applying our two-stage framework and selecting responses via $ODRP\textsubscript{HN}$, the quality of responses generated by smaller LLMs like $TinyLlama$ and $Llama2\text{-}7b$ surpasses that of $Llama2\text{-}70b$ in terms of response diversity and approaches the level of $gpt\text{-}3.5\text{-}turbo$. Regarding contextual coherence, $Llama2\text{-}13b$ see improvements that bring it in line with $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$, while $TinyLlama$ and $Llama2\text{-}7b$, although still trailing, narrow the gap significantly. Qualitatively, we note that responses selected by ODRPODRPHN are comparable to responses generated by $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$ in terms of naturalness and engagingness. These results underscore the effectiveness of our approach, enabling smaller LLMs to rival or exceed the capabilities of larger models, all while maintaining lower computational demands.
@@ 6.2 MRG Evaluation
Comparison with MI Response sets generated via MI tend to be semantically similar despite relatively high lexical diversity, as shown by low inter-response semantic scores and comparably higher lexical diversity scores in both automatic and human evaluations. This is likely due to the deterministic nature of logits during inference. Although sampling strategies (temperature scaling 10.5555/3305381.3305518 or nucleus sampling holtzman2020curiouscaseneuraltext) introduce stochasticity in decoding, generated logits remain deterministic, limiting semantic variation unless randomness is significantly increased, which could reduce contextual coherence.
@@ 7 Related Work
Prior work adopting a two-stage approach for open-domain dialogue typically involves generating multiple responses either through conditional generation based on pre-specified dialogue acts sakaeda-kawahara-2022-generate or by pooling outputs from variational and retrieval-based systems 10.1016/j.csl.2020.101071; Technology2021. However, these studies often focus on evaluating only the final selected response, without considering the diversity or contextual coherence of the entire set of generated responses. In contrast, our approach evaluates and optimizes the quality of the full set of responses, thereby enhancing the overall quality of the final selected response. Additionally, many of these methods have been applied to smaller language models, whereas to the best of our knowledge, our work is the first to introduce a two-stage generation framework LLMs. Other two-stage approaches broadly entail first generating a candidate response and instantiating it as the final response 10.1145/3551869, or generating a response in the first stage and further conditioning and refining the response in the second stage qian2024thinktwicehumanliketwostage; 9525043.
Regarding response selection, prior work has primarily concentrated on narrow criteria such as engagement sakaeda-kawahara-2022-generate, topical relevance 10.1016/j.csl.2020.101071; YUAN2024100087. Standard retrieval-based systems, in contrast, prioritize contextual coherence ijcai2021p627; su-etal-2024-dial. In our framework, we prioritize human preferences, considering a broader range of factors such as harmlessness and helpfulness, which are critical aspects for ensuring the real-world utility of response generation systems.
Dist-1 Dist-2 UE UniEval
TinyLlamaTinyLlama 0.16*/0.18* 0.51*/0.61* 0.21*/0.13* 0.74*/0.64*
- randrand 0.24*/0.20* 0.75*/0.70* 0.24/0.13* 0.76*/0.65*
- clscls 0.22*/0.25* 0.76*/0.74* 0.23/0.18* 0.78*/0.66*
- prefpref 0.25*/0.24* 0.73*/0.75* 0.24/0.18* 0.77*/0.70*
- ODRPODRP 0.28*/0.29 0.77/0.798 0.27/0.22* 0.81/0.72*
- ODRPODRPHN 0.31/0.31 0.79/0.82 0.30/0.26 0.83/0.76
$Llama2\text{-}7b$ 0.20*/0.22* 0.61*/0.69* 0.24*/0.21 0.83/0.72
- randrand 0.23*/0.30* 0.77*/0.78* 0.22*/0.19* 0.81*/0.69*
- clscls 0.30*/0.27* 0.79*/0.75* 0.23*/0.18* 0.80*/0.72
- prefpref 0.28/0.29* 0.77*/0.78* 0.24*/0.22 0.83/0.71*
- ODRPODRP 0.33/0.35 0.83/0.84 0.26/0.22 0.83/0.73
- ODRPODRPHN 0.35/0.36 0.83/0.85 0.29/0.24 0.85/0.73
$Llama2\text{-}13b$ 0.21*/0.23* 0.65*/0.72* 0.26*/0.24* 0.85/0.77*
- randrand 0.24*/0.28* 0.77*/0.76* 0.25*/0.24* 0.80*/0.72*
- clscls 0.30*/0.31* 0.80*/0.76* 0.29*/0.25 0.83*/0.77*
- prefpref 0.31/0.30* 0.79*/0.78 0.26*/0.29 0.82*/0.79
- ODRPODRP 0.33/0.34 0.85/0.79 0.32/0.30 0.85/0.81
- ODRPODRPHN 0.33/0.35 0.84/0.82 0.33/0.32 0.87/0.82
$Llama2\text{-}70b$ 0.31/0.32 0.72/0.80 0.28/0.26 0.86/0.79
$gpt\text{-}3.5\text{-}turbo$ 0.36/0.33 0.75/0.82 0.31/0.30 0.88/0.81
Win Tie Loss
TinyLlama ODRPODRPHN vs. TinyLlamaTinyLlama 85 9 6
ODRPODRPHN vs. randrand 76 16 18
ODRPODRPHN vs. clscls 60 29 11
ODRPODRPHN vs. prefpref 57 20 23
ODRPODRPHN vs. ODRPODRP 49 33 18
ODRPODRPHN vs. $Llama2\text{-}70b$ 30 35 35
ODRPODRPHN vs. $gpt\text{-}3.5\text{-}turbo$ 26 44 30
Llama2-7b ODRPODRPHN vs. $Llama2\text{-}7b$ 74 18 8
ODRPODRPHN vs. randrand 58 25 17
ODRPODRPHN vs. clscls 50 29 21
ODRPODRPHN vs. prefpref 47 27 23
ODRPODRPHN vs. ODRPODRP 46 30 24
ODRPODRPHN vs. $Llama2\text{-}70b$ 32 41 27
ODRPODRPHN vs. $gpt\text{-}3.5\text{-}turbo$ 28 48 24
Llama2-13b ODRPODRPHN vs. $Llama2\text{-}13b$ 50 33 17
ODRPODRPHN vs. randrand 51 24 25
ODRPODRPHN vs. clscls 44 34 22
ODRPODRPHN vs. prefpref 42 30 28
ODRPODRPHN vs. ODRPODRP 41 32 27
ODRPODRPHN vs. $Llama2\text{-}70b$ 38 39 23
ODRPODRPHN vs. $gpt\text{-}3.5\text{-}turbo$ 37 40 23